Starting at Joe's Gas Station, a car travels due north on Avenue A for 3.4 miles, then turns right and travels on Cedar Street for 1.0 miles, and then turns right again and travels on Washington Avenue for 0.6 miles. How far is the car from Joe's Gas Station?
Answer by stanbon(48558)

Starting at Joe's Gas Station, a car travels due north on Avenue A for 3.4 miles, then turns right and travels on Cedar Street for 1.0 miles, and then turns right again and travels on Washington Avenue for 0.6 miles. How far is the car from Joe's Gas Station?
-----------------------------------------------
North travel: 3.4 mi
East travel: 1 mi
South travel 0.6 mi
-------------------------
Total change: East 1 mi and (3.4-0.6) = 2.8 North
---------------------
Use Pythagoras:
distance^2 = 1^2 + 2.8^2
d^2 = 1 + 7.84
d = 2.97 miles (distance from starting point to ending point)
===================
Cheers,
Stan H.